OpenStudy (anonymous):

How do you find the X an Y axis of a quadratic function? I need to know how to do this WITHOUT graphing :) You can use: y = x2 + 2x - 3? That quadratic equation as an example, if you want.

OpenStudy (saifoo.khan):

when x intercpts y = 0 when y intercpts x = 0

OpenStudy (anonymous):

Couldyou explain that a little more? Sorry, I don't quite understand what you mean, haha.

OpenStudy (saifoo.khan):

Sure, No Problem! =)

OpenStudy (anonymous):

Awesome ^^ Thank you!

OpenStudy (saifoo.khan):

when x intercpts y = 0 y = x2 + 2x - 3 0 = x^2 + 2x -3 x = 1, x=-3 when y intercpts x = 0 y = x2 + 2x - 3 y = (0)^2 + 2(0) - 3 y = -3

OpenStudy (anonymous):

OH! I see you what you mean. Thanks! That's super easy :]

OpenStudy (saifoo.khan):

Whenever a line cuts the x axis, y = 0 Whenever a line cuts the y axis, x = 0.

OpenStudy (saifoo.khan):

Surely it is! xD